6. A ball is tied at the end of a rope and swings in a circle at a constant speed, doing 60 revolutions a minute. What is the frequency and which period does this movement have?
1
Expert's answer
2020-09-21T08:28:52-0400

The frequency of rotation

"f=\frac{N}{t}=\rm\frac{60\: rev}{1\:min}=1\: rev/s"

The period of rotation

"T=\frac{1}{f}=\rm \frac{1}{1\:rev/s}=1\:s"